# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

### v1

@@ Abstract
We present MoGe, a powerful model for recovering 3D geometry from monocular open-domain images. Given a single image, our model directly predicts a 3D point map of the captured scene with an affine-invariant representation, which is agnostic to true global scale and shift. This new representation precludes ambiguous supervision in training and facilitate effective geometry learning. Furthermore, we propose a set of novel global and local geometry supervisions that empower the model to learn high-quality geometry. These include a robust, optimal, and efficient point cloud alignment solver for accurate global shape learning, and a multi-scale local geometry loss promoting precise local geometry supervision. We train our model on a large, mixed dataset and demonstrate its strong generalizability and high accuracy. In our comprehensive evaluation on diverse unseen datasets, our model significantly outperforms state-of-the-art methods across all tasks, including monocular estimation of 3D point map, depth map, and camera field of view. Code and models will be released on our project page.
@@
Figure 1: Given any image, our method reconstructs an affine-invariant 3D point map of the scene and can also produce a depth map and the camera focal length. The model yields high-quality shapes and generalizes well to open-domain images. (Best view with zoom)
@@ 1 Introduction
For monocular geometry estimation (MGE), a common approach involves first estimating a depth map, subject to an unknown scale (and shift), and then combining it with camera intrinsics to recover the 3D shape via unprojection [69, 41]. For the former, significant advancements in monocular depth estimation (MDE) have been achieved in recent years by training on large-scale data [71, 65, 29, 20]. However, inferring camera intrinsics (e.g., focal length) from single images remains challenging due to the high degree of ambiguity when strong geometric cues are absent. Inaccurate camera parameters can lead to significant geometry distortions even when used with ground truth depth maps. Recently, DUSt3R [57] proposed predicting a 3D point map from an image by mapping each pixel to a free 3D point, thereby bypassing camera parameter estimation during geometry recovery. Although the model can estimate monocular geometry, it is primarily designed for multi-view scenarios and is trained on image pair inputs.
@@
More importantly, our findings indicate that the design of training supervisions is crucial to achieve robust and accurate point map prediction from single images. Similar to previous MDE approaches, we need to compute a global scaling factor and a translation to align the predicted results with the ground truth during training. However, previous methods for this global alignment calculation are either sensitive to outliers or solved with coarse approximations, leading to unsatisfactory supervision. We propose a Robust, Optimal, and Efficient (ROE) global alignment solver to resolve scale and shift for the affine-invariant pointmap loss, which substantially improved the training effectiveness and final accuracy. On the other hand, the effective learning of local, region-wise geometry has been largely neglected previously. In monocular geometry estimation, the relative distance between different objects can be ambiguous, which hinders precise local geometry learning when a global alignment computed with all objects is applied. In light of this, we propose a multi-scale local geometry loss, which imposes penalties on the local discrepancies of 3D point clouds, each under independent and optimal affine alignment. This design significantly boosted the accuracy of our local geometry prediction.
We train our model on a large-scale training corpus sourced from various existing datasets. Our model demonstrates strong generalizability and accuracy in monocular geometry estimation for open-domain images (some examples are shown in Fig. 1). The zero-shot evaluation on eight unseen datasets shows that it outperforms previous MGE methods that output point clouds by a large margin (over 35% error reduction compared to the previous best). Moreover, our method also significantly outperforms previous approaches that focus on the sub-tasks of MDE and camera FoV (20$\sim$30error reduction for the former and over 20% for the latter). Figure 2 shows the performance ranking of different methods in our evaluation, where our method ranks top across all tasks/metrics and among all compared methods.
Our contributions are summarized as follows: 1) We propose a new direct MGE method for open-domain images using affine-invariant point maps. 2) We establish a set of novel and effective global and local supervisions for robust and precise geometry recovery. 3) We demonstrate the substantial enhancement in performance of our model over existing methods on MGE, MDE, and camera FOV estimation across diverse datasets. We hope that our method can serve as a strong and versatile foundational model for solving monocular geometry problems, facilitating applications such as 3D-aware image editing, image relighting, depth-to-image synthesis, novel view synthesis, and 3D scene understanding, as well as providing initial geometrical priors for video-based or multiview-based 3D reconstruction.
@@ 2 Related Work
Compared to depth map, 3D point map offers a more straightforward representation of 3D geometry. Monocular point map estimation aims to recover free 3D points for each pixel on a single image. Several approaches tackle this challenge by predicting camera parameters alongside depth estimation. For example, LeReS [69, 70] introduced a two-stage pipeline with an affine-invariant depth prediction module followed by a point cloud module for recovering shift and camera focal length. UniDepth [41] proposed a self-promptable camera module that predicts dense camera representation to condition the subsequent depth estimation module. DUSt3R [57] employs a more straightforward end-to-end model to directly map two-view images to camera space point maps, and it can adapt to monocular scenarios by using two identical images as input. It uses the camera-space scale-invariant pointmap which incorporates camera shift information. This representation can suffer from the focal-distance ambiguity and jeopardize pointmap learning. In contrast, our approach is designed to handle monocular input and we employ affine-invariant point map with meticulously designed training supervisions for more effective geometry learning.
@@
Estimating camera intrinsics is a fundamental task in 3D vision. While traditional multiview camera calibration using special patterns is well studied [73], estimating camera intrinsics from single images is ill-posed and remains challenging. For the latter, earlier works utilized known 3D shapes [59] or vanishing points [13] to solve the problem. More recently, learning-based approaches have been proposed [61, 7, 32, 28, 75] to handle in-the-wild images, but the results are far from satisfactory. In this work, we demonstrate that our camera parameter estimation derived from point maps can achieve state-of-the-art accuracy and offer remarkable generalizability.
@@ 3.1 Affine-invariant point map
The affine-invariant point map representation can be used to recover both camera shift and focal length. Given the predicted 3D points $(x_{i},y_{i},z_{i})$ and their corresponding 2D pixels $(u_{i},v_{i})$, we solve for camera focal length $f$ and $Z$-axis shift by minimizing the projection error:
@@
where $t_{z}^{\prime}=t_{z}/s$. Eq. 2 can be efficiently solved with a fast iterative optimization algorithm, which converges within 10 iterations, taking approximately $3$ms. More details about the solution can be found in the suppl. materials.
@@ 3.2 Training Objectives
$\mathcal{L}_{G}=\sum_{i\in{\mathcal{M}}}{\frac{1}{{z}_{i}}\|s\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\|_{1}},$ (3)
@@
To apply the global loss $\mathcal{L}_{G}$ for training, we need to first determine $s$ and $\mathbf{t}$. Previous methods for affine-invariant depth estimation often use coarse approximations for this alignment. For example, a predominate strategy is to use the median depth values of the two sets as the anchor to compute shift followed by scaling factor calculation [42, 65, 66]. The range normalization strategy used by [29, 20] can be viewed as computing shift using minimal or a predefined fraction of depth values. All these simple strategies are non-optimal for the problem setup and could lead to unsatisfactory supervision.
In this work, we propose an optimal solver to the alignment parameters. Specifically, we determine $s^{*}$ and $\mathbf{t}^{*}$ via
@@
where $t_{x}$ and $t_{y}$ are 00. One possible approach to solve this equation is to formulate it as an absolute residuals optimization problem and solve it using linear programming techniques such as simplex algorithms and the interior-point method [8]. However, these methods are known to have high computational complexity, often exceeding $O(N^{3})$. Given thousands of points, a typical implementation may require computation time of a few seconds or even higher, which is inefficient for network training.
Instead, we develop an efficient parallelized searching algorithm to solve it optimally. Our theoretical analysis (Sec. A.2) demonstrates that the minimum must occur when $s\hat{z}^{(k)}+t_{z}-z^{(k)}=0$ for some $k=1,\ldots,N,\ k\in\mathcal{M}$. This insight allows us to substitute $t_{z}$ with $s$ and decompose the problem into a series of one-dimensional subproblems with scale variable only, which can be solved in parallel. The overall computational complexity is $O(N^{2}\log N)$ and efficient training is achieved on GPU. A high-level description of this solution is presented in Algorithm 1 and more details can be found in the suppl. materials.
In practice, we noticed that Eq. 4 is still sensitive to outliers even if the $L_{1}$ error norm has been used. If the model wrongly predicts a close foreground edge pixel as a far background point, the objective will be dominated by the incorrect edge prediction. To further improve the robustness, we apply a clipping operator $\operatorname{clip}\left(\cdot,\tau\right)=\max\big{(}\min(\cdot,\tau),-\tau\big{)}$ on the per-point residuals, which brings accuracy gains as we will show in the experiments. Adding the clipping operator makes the problem non-convex, but our optimal solver is still applicable with minor modifications.
@@
Algorithm 1 Optimal alignment solver to Eq. 4
function solve_subproblem
    Enumerating non-differentiable points of objective function,
         at zero per-coordinate alignment residuals.
    Determining extrema by derivative computation.
@@
for index $k=1,...,N$ do $\triangleright$ parallel computation
    Formulate subproblem by substituting $t_{z}$ with $z^{(k)}\!-\!s\hat{z}^{(k)}$.
    Solve scale $s^{(k)}$ and $l^{(k)}$ via SOLVE_SUBPROBLEM.
    Obtain translation $t_{z}^{(k)}$ as $z^{(k)}\!-\!s^{(k)}\hat{z}^{(k)}$.
@@
where $r_{j}$ is the radius. We set $r_{j}=\alpha\cdot z_{j}\cdot\frac{\sqrt{W^{2}+H^{2}}}{2\cdot f}$ where $z_{j}$ is the $z$-coordinate of $p_{j}$, $f$ is the focal length, and $W$ and $H$ are image width and height. This way, the hyper-parameter $\alpha\in(0,1)$ approximates the proportion of the projected sphere’s diameter to the image diagonal.
Then, we apply the aforementioned ROE alignment solver to align the point maps and compute error. For each sphere scale parameter $\alpha$, we sample a set of anchor points $\mathcal{H}_{\alpha}$ and compute the loss as
$\mathcal{L}_{S(\alpha)}=\sum_{j\in\mathcal{H}_{\alpha}}l_{{\mathcal{S}}_{j}}=\sum_{j\in\mathcal{H}_{\alpha}}\sum_{i\in{S}_{i}}\frac{1}{z_{i}}\left\|s^{*}\hat{\mathbf{p}}_{i}+\mathbf{t}^{*}-{\mathbf{p}}_{i}\right\|_{1}$ (6)
In practice, we use three $\alpha$ scales ranging from coarse to fine: $1/4$, $1/16$, and $1/64$, denoted as $\mathcal{L}_{S_{1}}$, $\mathcal{L}_{S_{2}}$ and $\mathcal{L}_{S_{3}}$, respectively.
@@
$\mathcal{L}_{N}=\sum_{i\in{\mathcal{M}}}\angle(\hat{\mathbf{n}}_{i},\mathbf{n}_{i})$ (7)
where the normal $\mathbf{n}_{i}$ of pixel $i$ is obtained from the cross product of its adjacent edges on the image grid.
@@
The distance to infinity region (e.g., sky) in outdoor scenes and plain backgrounds in object-only images are undefined. We add a single-channel head to our model for predicting the mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ where the distance is valid:
$\mathcal{L}_{M}=\|\hat{\mathbf{M}}-(1-{\mathbf{M}}_{\text{inf}})\|_{2}^{2}$ (8)
where ${\mathbf{M}}_{\text{inf}}$ is the infinity mask label. For synthetic data, the ground truth masks are readily available, while for real outdoor scenes, we use SegFormer [62] to obtain sky masks.
At inference time, we use a threshold of $0.5$ to binarize the mask prediction. Recovery camera focal and shift involves masked points only, excluding the regions of meaningless geometry.
@@ 3.3 Training data
Recently, numerous synthetic and real RGBD datasets have been utilized in large-scale training for monocular geometry estimation. In this work, we collected $21$ datasets from various domains to train our model. The total number of collected frames is around $9$ million. Details of these datasets and their mixing ratios can be found in the suppl. materials.
Since the quality of data labels are varied, we apply different loss function combinations tailored for different data sources. As shown in Table 1, higher quality and less noisy data are assigned with finer-level loss functions. To address imperfections and potential outliers in real labels, we exclude the top $5\%$ of per-pixel loss values for real data. Following DINOv2 [40], we rebalance the collected datasets using image retrieval from a curated image set to address uneven or biased data distributions. See the suppl. materials for details.
Table 1: Label quality and applied losses for different training sources.
Source Label quality Applied losses
Synthetic Perfect $\infty$ Dense
SfM/MV Recon High $\infty$ Dense&Partial
LiDAR/Laser High $\sim 100$m Sparse
Kinect Medium $\sim 10$m Dense
@@ 4 Experiments
Scale-invariant depth
@@
Affine-invariant depth
@@
Affine-invariant disparity
@@
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
Table 3: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking.
@@
We employ a ViT [15] encoder and a light-weight CNN-based upsampler as the decoder. The encoder is pretrained with DINOv2 [40]. The initial learning rates for the encoder and decoder are set to $5\times 10^{-6}$ and $5\times 10^{-5}$, with a decay factor of $5$ applied for every 100K iterations. The training batch size is $256$. Our model is trained on images with varying aspect ratio and resolutions, where aspect ratios are uniformly sampled between $1:2$ and $2:1$, and pixel counts range from 250K to 500K. We apply several image augmentations, including color jittering, Gaussian blurring, JPEG compression-decompression, and random cropping. The principal point of the cropped image is aligned with its center using a perspective crop.

### v2

@@ Abstract
We present MoGe, a powerful model for recovering 3D geometry from monocular open-domain images. Given a single image, our model directly predicts a 3D point map of the captured scene with an affine-invariant representation, which is agnostic to true global scale and shift. This new representation precludes ambiguous supervision in training and facilitates effective geometry learning. Furthermore, we propose a set of novel global and local geometry supervision techniques that empower the model to learn high-quality geometry. These include a robust, optimal, and efficient point cloud alignment solver for accurate global shape learning, and a multi-scale local geometry loss promoting precise local geometry supervision. We train our model on a large, mixed dataset and demonstrate its strong generalizability and high accuracy. In our comprehensive evaluation on diverse unseen datasets, our model significantly outperforms state-of-the-art methods across all tasks, including monocular estimation of 3D point map, depth map, and camera field of view. Code and models can be found on our project page.
@@
Figure 1: Given any image, our method reconstructs an affine-invariant 3D point map of the scene and can also produce a depth map and the camera focal length. The model yields high-quality shapes and generalizes well to open-domain images. (Best viewed with zoom)
@@ 1 Introduction
For monocular geometry estimation (MGE), a common approach involves first estimating a depth map, subject to an unknown scale (and shift), and then combining it with camera intrinsics to recover the 3D shape via unprojection [71, 42]. For the former, significant advancements in monocular depth estimation (MDE) have been achieved in recent years by training on large-scale data [73, 26, 67, 68, 43] or leveraging powerful pretrained models [29, 20, 23]. However, inferring camera intrinsics (e.g., focal length) from single images remains challenging due to the high degree of ambiguity when strong geometric cues are absent. Inaccurate camera parameters can lead to significant geometry distortions even when used with ground truth depth maps. Recently, DUSt3R [59] proposed predicting a 3D point map from an image by mapping each pixel to a free 3D point, thereby bypassing camera parameter estimation during geometry recovery. Although the model can estimate monocular geometry, it is primarily designed for multi-view scenarios and is trained on image pair inputs.
@@
More importantly, our findings indicate that the design of training supervisions is crucial to achieve robust and accurate point map prediction from single images. Similar to previous MDE approaches, a global scaling factor and translation are required to align predictions with ground truth during training. However, existing methods for this global alignment calculation are either sensitive to outliers or solved with coarse approximations, leading to suboptimal supervision. We propose a Robust, Optimal, and Efficient (ROE) global alignment solver to resolve scale and shift for the affine-invariant point map loss, which substantially improves the training effectiveness and final accuracy. On the other hand, the effective learning of local, region-specific geometry has been largely neglected previously. In monocular geometry estimation, the relative distances between different objects can be ambiguous, which hinders precise local geometry learning when only a global alignment is applied. In light of this, we propose a multi-scale local geometry loss, which penalizes local discrepancies in 3D point clouds under independent affine alignments. This design significantly boosted the accuracy of our local geometry prediction.
We train our model on large-scale data sourced from various existing datasets. Our model demonstrates strong generalizability and accuracy in monocular geometry estimation for open-domain images (examples in Fig. 1). Zero-shot evaluation on eight unseen datasets shows a significant improvement over previous MGE methods that output point clouds (over 35% error reduction compared to the previous best). Moreover, our method significantly outperforms prior approaches that focus on the sub-tasks of MDE and camera FOV (20$\sim$30error reduction for the former and over 20% for the latter). Figure 2 shows that our method ranks top across all tasks and metrics in the evaluation.
Our contributions are summarized as follows: 1) We propose a new direct MGE method for open-domain images using affine-invariant point maps. 2) We establish a set of novel and effective global and local supervisions for robust and precise geometry recovery. 3) We demonstrate the substantial enhancement in performance of our model over existing methods on MGE, MDE, and camera FOV estimation across diverse datasets. We hope that our method serves as a versatile foundational model for monocular geometry tasks, facilitating applications such as 3D-aware image editing, depth-to-image synthesis, novel view synthesis, and 3D scene understanding, as well as providing initial geometrical priors for video or multiview-based 3D reconstruction.
@@ 2 Related Work
Compared to depth maps, 3D point map offers a more straightforward representation of 3D geometry. Monocular point map estimation aims to recover free 3D points for each pixel in a single image. Several approaches address this by predicting camera parameters alongside depth. For example, LeReS [71, 72] introduced a two-stage pipeline with affine-invariant depth prediction followed by a point cloud module to recover shift and camera focal length. UniDepth [42] proposed a self-promptable camera module that predicts dense camera representation to condition the subsequent depth estimation module. DUSt3R [59] employs an end-to-end model to directly map two-view images to camera space point maps, adaptable to monocular scenarios by using two identical input images. However, its scale-invariant point map can suffer from the focal-distance ambiguity. In contrast, our approach is designed to handle monocular input and we employ affine-invariant point map with meticulously designed training supervisions for more effective geometry learning.
@@
Estimating camera intrinsics is a fundamental task in 3D vision. While traditional multiview camera calibration using special patterns is well studied [75], estimating camera intrinsics from single images is ill-posed and remains challenging. For the latter, earlier works utilized known 3D shapes [61] or vanishing points [13]. More recently, learning-based approaches have been explored [63, 7, 33, 28, 77] to handle in-the-wild images, but the results are far from satisfactory. In this work, we demonstrate that our camera parameter estimation derived from point maps can achieve state-of-the-art accuracy and offer remarkable generalizability.
@@ 3.1 Affine-invariant point map
The affine-invariant point map representation can be used to recover both camera shift and focal length. Given the predicted 3D points $(x_{i},y_{i},z_{i})$ and their corresponding 2D pixels $(u_{i},v_{i})$, we solve for camera focal length prediction $f$ and $Z$-axis shift $t_{z}^{\prime}$ by minimizing the projection error:
@@
where $t_{z}^{\prime}=t_{z}/s$. Eq. 2 can be efficiently solved with a fast iterative optimization algorithm, which typically converges within 10 iterations, taking approximately $3$ms. More details about the solution can be found in the suppl. materials.
@@ 3.2 Training Objectives
$\mathcal{L}_{G}=\sum_{i\in{\mathcal{M}}}{\frac{1}{{z}_{i}}\|s\hat{\mathbf{p}}_{i}+\mathbf{t}-{\mathbf{p}}_{i}\|_{1}},$ (3)
@@
To apply the global loss $\mathcal{L}_{G}$ for training, we need to first determine $s$ and $\mathbf{t}$. Previous affine-invariant depth estimation methods often use rough alignment approximations, like using the median depth of two sets as an anchor to compute shift, followed by scaling [43, 67, 68], or normalizing depth range [29, 20]. These simple strategies yield suboptimal alignment and could lead to unsatisfactory supervision.
In this work, we propose a solver to the optimal alignment parameters. Specifically, we determine $s^{*}$ and $\mathbf{t}^{*}$ via
@@
where $t_{x}$ and $t_{y}$ are 00. One approach to solve this equation is to frame it as an absolute residuals optimization problem and use linear programming methods such as simplex algorithms and the interior-point method [8]. However, these methods are known to have high computational complexity, often exceeding $O(N^{3})$, making them inefficient for network training with thousands of points, where computation can take seconds. Instead, we develop an efficient parallelized searching algorithm. With the fact that the optimum must occur when $s\hat{z}_{k}+t_{z}=z_{k}$ for some index $k$, we replace $t_{z}$ with $s$ and break the problem into a series of parallel one-dimensional subproblems. This reduces the complexity to $O(N^{2}\log N)$ and enables efficient GPU-based training. A high-level description of this solution is presented in Algorithm 1 with more details in the suppl. materials.
In practice, we found that Eq. 4 is still sensitive to outliers occasionally even with the $L_{1}$ error norm. If the model mistakenly predicts a close foreground edge pixel as a distant background point, the objective would be dominated by this incorrect prediction. To further improve the robustness, we apply truncation $\min(\cdot,\tau)$ to the absolute residuals. While this truncation introduces non-convexity, our solver remains applicable with minor adjustments.
@@
Algorithm 1 Overview of ROE alignment
function SolveSubproblem()
    Enumerate breakpoints of the piecewise linear function.
    Find extrema among the breakpoints by their derivatives.
@@
for index $k=1$ to $N$ do $\triangleright$ parallel computation
    Formulate subproblem by substituting $t_{z}$ with $z_{k}\!-s^{(k)}\hat{z}_{k}$.
    Solve scale $s^{(k)}$ and $l^{(k)}$ via SolveSubproblem().
    Obtain translation $t_{z}^{(k)}$ as $z_{k}\!-\!s^{(k)}\hat{z}_{k}$.
@@
where $r_{j}$ is the radius. We set $r_{j}=\alpha\cdot z_{j}\cdot\frac{\sqrt{W^{2}+H^{2}}}{2\cdot f}$ where $z_{j}$ is the $z$-coordinate of $p_{j}$, $f$ is the ground truth focal length, and $W$ and $H$ are image width and height. This way, the hyper-parameter $\alpha\in(0,1)$ approximates the proportion of the projected sphere’s diameter to the image diagonal.
Then, we apply the aforementioned ROE alignment solver to align the point maps with optimal $(s_{j}^{*},\mathbf{t}_{j}^{*})$ and compute error. For each sphere scale parameter $\alpha$, we sample a set of anchor points $\mathcal{H}_{\alpha}$ and compute the loss as
$\mathcal{L}_{S(\alpha)}=\!\sum_{j\in\mathcal{H}_{\alpha}}l_{{\mathcal{S}}_{j}}\!=\!\sum_{j\in\mathcal{H}_{\alpha}}\!\sum_{i\in{S}_{j}}\frac{1}{z_{i}}\left\|s_{j}^{*}\hat{\mathbf{p}}_{i}+\mathbf{t}_{j}^{*}-{\mathbf{p}}_{i}\right\|_{1}.$ (6)
In practice, we use three $\alpha$ scales from coarse to fine: $1/4$, $1/16$, and $1/64$, denoted as $\mathcal{L}_{S_{1}}$, $\mathcal{L}_{S_{2}}$ and $\mathcal{L}_{S_{3}}$, respectively.
@@
$\mathcal{L}_{N}=\sum_{i\in{\mathcal{M}}}\angle(\hat{\mathbf{n}}_{i},\mathbf{n}_{i}),$ (7)
where the normal $\hat{\mathbf{n}}_{i}$ of pixel $i$ is obtained from the cross product of its adjacent edges on the image grid, $\mathbf{n}_{i}$ is the ground truth, and (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ) measures the angle difference between the two vectors.
@@
The infinity region (e.g., sky) in outdoor scenes and plain backgrounds in object-only images have undefined geometry. We use a single-channel output head for predicting the mask $\hat{\mathbf{M}}\in\mathbb{R}^{H\times W}$ of valid points:
$\mathcal{L}_{M}=\|\hat{\mathbf{M}}-(1-{\mathbf{M}}_{\text{inf}})\|_{2}^{2},$ (8)
where ${\mathbf{M}}_{\text{inf}}$ is the infinity mask label. For synthetic data, the ground truth masks are readily available, while for real outdoor scenes, we use SegFormer [64] to obtain sky masks. At inference time, the predicted mask is binarized with a threshold of $0.5$. The recovery of focal and shift involves masked points, excluding the regions of invalid geometry.
@@ 3.3 Training data
We collected 21 publicly available datasets from various domains [21, 62, 3, 14, 69, 35, 74, 51, 57, 45, 58, 40, 34, 19, 27, 37, 76, 47, 60, 25, 12], totaling around 9 million frames for training our model. Since the quality of data labels are varied, we apply tailored loss function combinations for different data sources. Higher-quality and less noisy data are assigned with finer-level loss functions. To address imperfections and potential outliers in real labels, we exclude the top $5\%$ of per-pixel loss values for real data. Following DINOv2 [41], we balance the weights of collected datasets via image retrieval from a curated image set. Details of training datasets can be found in the suppl. materials.
@@ 4 Experiments
     Scale-invariant depth map
@@
     Affine-invariant depth map
@@
Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.00
     Affine-invariant disparity map
@@
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
Table 2: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking. An extensive comparison is found in the suppl. materials.
@@
We use a ViT [15] encoder pretrained with DINOv2 [41] and a lightweight CNN-based upsampler as the decoder. The encoder and decoder have initial learning rates of $5\times 10^{-6}$ and $5\times 10^{-5}$, with a decay factor of 5 every 100K iterations. The training batch size is $256$. Our model is trained on images with aspect ratios uniformly sampled between $1:2$ and $2:1$, and pixel counts from 250K to 500K. We apply several image augmentations, including color jittering, Gaussian blurring, JPEG compression-decompression, and random cropping. The principal point of the cropped image is aligned with its center using a perspective crop.